OpenStudy (anonymous):

how do I find sin 5ᴨ/8, cos 5ᴨ/8 and tan 5ᴨ/8 using trig identities

OpenStudy (anonymous):

What is ᴨ

OpenStudy (anonymous):

use formula for sin4x

OpenStudy (anonymous):

the problem doesn't say...just says to use trigonometric identities to find the values

OpenStudy (anonymous):

But what is "ᴨ" is it theta or something else. What is it?

OpenStudy (anonymous):

expand sin4x, put x as 5pi/8 and equate the expa nsion of sin4x to 1

OpenStudy (anonymous):

Is it pi?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sin ((270-45)/2)=sin (5(pi)/8)

OpenStudy (anonymous):

Do you understand this step?

OpenStudy (anonymous):

Do you understand the step "Sin ((270-45)/2)=sin (5(pi)/8)"

OpenStudy (anonymous):

not really, but i think you got the 270 from 3pi/2 and the 45 from pi/4....

OpenStudy (anonymous):

Well, I am trying to express the thing in degree format. (Sorry for using jumbled form). By the term I mean 180 degree

OpenStudy (anonymous):

By the term pi mean 180 degree

OpenStudy (anonymous):

\[Sin ((3\pi/2-\pi/4)/2)=sin (5\pi/8)\]

OpenStudy (anonymous):

Now does that makes sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

*make

OpenStudy (anonymous):

Now can you do the rest, or I will have to show you every steps.

OpenStudy (anonymous):

so, sin (225/2)=sin (112.5)? you might have to show me the steps.

OpenStudy (anonymous):

or is it sin (5pi/4 divided by 2)

OpenStudy (anonymous):

Yes, thats true, but thats not helpful here. \[Sin ((3\pi/2-\pi/4)/2)=sin (5\pi/8)\] \[=>sin (5\pi/8)=\sqrt(0.5(1-cos(3\pi/2-\pi/4))\]

OpenStudy (anonymous):

Now you got to use addition formula for cos to solve the part with cos, and then it will be done

OpenStudy (anonymous):

I am totally lost .....this does not look at all familiar.....

OpenStudy (anonymous):

are you familiar with double angle formula for trig?

OpenStudy (anonymous):

i'm sure we went over them....but nothing coming to mind

OpenStudy (anonymous):

Then you can't expect to get this thing within your understandability range

OpenStudy (anonymous):

that's why I am having trouble....thought if someone walked me through it, it would make sense

OpenStudy (anonymous):

If you wish I can tutor you (in exchange of some nominal payment) :)

OpenStudy (anonymous):

Seems like you are not interested. So may I leave?

OpenStudy (anonymous):

didn't know i had to pay for help

OpenStudy (anonymous):

No no not that

OpenStudy (anonymous):

you can get free help here for ever

OpenStudy (anonymous):

But I don't think they will be very helpful

OpenStudy (anonymous):

So incase you need personal help

OpenStudy (anonymous):

I mean through audio chatting and virtual black board

OpenStudy (anonymous):

I can't teach you everything here. This place is too inconvenient to type even a simple equation

OpenStudy (anonymous):

I mean no one can

OpenStudy (anonymous):

thanks for the offer, but rather just stay here

OpenStudy (anonymous):

Thats great. No problem